A boat travels 20 km/h in still water. If this boat goes upstream in a river flowing at a rate of 6 km/h, for 10 km, and then returns back downstream to its starting point, what is the average speed for the whole trip in km/h?
Answer by mananth(16946)

A boat travels 20 km/h in still water. If this boat goes upstream in a river flowing at a rate of 6 km/h, for 10 km, and then returns back downstream to its starting point, what is the average speed for the whole trip in km/h?
Average speed = Total distance traveled / total time taken
Distance one way 10 km
boat travels 20 km/h in still water
speed of current = 6 km/h
upstream speed = 29-6 =14km/h
upstream time = 10/14-5/7
downstream speed = 26km/h
downstream time = 10/26= 5/13
Total time taken = 5/7 + 5/13= (65+35)/91 = 100/91
Speed = 20/(100/91)=18.2
the correct average speed for the whole trip is 18.2 km/h.
=